OpenStudy (anonymous):

What are the solutions of 3x2 + 15 = -6x? -1 ± i -1 ± 2i -2 ± i -2 ± 2i

OpenStudy (anonymous):

@student_basil?

OpenStudy (anonymous):

@amistre64?

OpenStudy (anonymous):

\[ x= -1\pm 2 i \]

OpenStudy (anonymous):

\[ 3 x^2+15=-6 x \\ 3 x^2 + 6 x +15=0\\ b^2-4 a c=36 -180= - 1 44\\ \sqrt{b^2-4 a c}= 12 i\\ x=\frac{-6\pm 12i}{6}= -1\pm 2i \]